OpenStudy (anonymous):

If a airplane flew at an average speed of 432 miles per hour for 2 and 1/2 hours would the plane have flew 1080???

OpenStudy (anonymous):

MILES! 1080 miles

OpenStudy (anonymous):

2 and 1/2=5/2 hours (5/2)*(432)=1080 yes you r correct answer good cluck your home work

OpenStudy (anonymous):

XD thanks